The U.S. population in 1990 was approximately 250 million, and the average growth rate for the past 30 years gives a doubling time of 66 years.
using the formula P(in millions)=250*2^(y-1990)/66
What will the population of the United States be in 2025 if this growth rate continues?
I am so stumped on this problem!

Answer by fractalier(2101)

All they're asking you to do is to substitute 2025 in for y and find the population P...you'll need your calculator though...so we get
P(in millions) = 250 * 2^[(y-1990)/66]
P(in millions) = 250 * 2^[(2025 - 1990) / 66]
P(in millions) = 250 * 2^[35 / 66]
P(in millions) = 361